OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (−6, −6) and (−3, 1).

OpenStudy (mimi_x3):

Use the gradient formula: \[m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}} \]

OpenStudy (mertsj):

Remember, subtract the y values. The subtract the x values. Put the first number over the second. rise over run (-6-1) over (-6-(-3))=-7/-3=7/3

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

Rise/Run will give you the equation of the line or you can use y-y1=m(x-x1)